The equity sections from Atticus Group’s 2013 and 2014 year-end balance sheets follow.
  Stockholders’ Equity (December 31, 2013)
  Common stock—$4 par value, 100,000 shares
    authorized, 40,000 shares issued and outstanding
$ 160,000
   Paid-in capital in excess of par value, common stock 120,000
   Retained earnings 320,000
    
   Total stockholders’ equity $ 600,000
    

         

  Stockholders’ Equity (December 31, 2014)
  Common stock—$4 par value, 100,000 shares
    authorized, 47,400 shares issued, 3,000 shares in treasury
$ 189,600
  Paid-in capital in excess of par value, common stock 179,200
  Retained earnings ($30,000 restricted by treasury stock) 400,000
   
   768,800
  Less cost of treasury stock (30,000 )
   
  Total stockholders’ equity $ 738,800
    

   

The following transactions and events affected its equity during year 2014.
     
Jan. 5 Declared a $0.50 per share cash dividend, date of record January 10.
Mar. 20 Purchased treasury stock for cash.
Apr. 5 Declared a $0.50 per share cash dividend, date of record April 10.
July 5 Declared a $0.50 per share cash dividend, date of record July 10.
July 31 Declared a 20% stock dividend when the stock’s market value is $12 per share.
Aug. 14 Issued the stock dividend that was declared on July 31.
Oct. 5

Declared a $0.50 per share cash dividend, date of record October 10.

How many common shares are outstanding on each cash dividend date?
What is the total dollar amount for each of the four cash dividends?

What is the amount of the capitalization of retained earnings for the stock dividend?

What is the per share cost of the treasury stock purchased?

How much net income did the company earn during year 2014?
Answer #1
Concepts and reason

Financial statements: These statements are reported by company at the end of each accounting year. They provide details about a company's annual earnings, cash available at the end of the year, assets and liabilities, stockholders' equity, and retained earnings. The basic financial statements are income statement, balance sheet, and statement of cash flows.

Balance sheet: This statement provides details of available assets and owed liabilities from investing and financial activities of the business. This statement reveals the financial health of company. So, this statement is also called as Statement of Financial Position. The users could estimate whether the company has enough assets to pay off its liabilities. The main components of balance sheet are assets, liabilities, and stockholders' equity.

Stockholders’ equity: Stockholders’ equity refers to the shareholders claims on the assets or resources of a company, and so known also as net assets of the company, which are assets minus liabilities. Examples: Retained Earnings, Dividends, and Capital.

Preferred stock: Preferred stock is the stock issued by the company indicating that the holders of the preferred stock are given priority while issuing dividend. They are completely settled and in case of arrears the dividend or any amount is settled in future year.

Paid in capital: Paid-in capital is the part of stockholders ‘equity appears complete the purchase of the company’s stock by investors. Paid-in capital refers to the capital stock. The main source of paid-in capital is capital contributed by investors or stockholders in the company.

Fundamentals

Common stock: It is a form of corporate equity ownership. It is one of the forms of securities that it is issued to common shareholders. They are entitled to dividend and repayment of capital after payment is made to preference shareholders. It will be reported in balance sheet in liabilities side.

Authorized shares: Authorized shares are the number of shares granted by state charter and are mentioned in its articles of association that are to be issued by company.

Outstanding shares: Outstanding shares are the number of shares that are issued by company and held by stockholders.

Issued shares: Issued shares are the number of shares that are allotted to stockholders out of the legally authorized shares

Treasury stocks: Treasury stocks are the company’s own stocks reacquired by the company which were issued previously. These are considered as issued stock but do not carry outstanding status. Thus, it does not have any voting right or dividend rights.

Dividends: The cash or stock distributed by the corporation to the investors, in return for the amount invested by them in the corporation is known as dividends.

Retained earnings: The amount of earnings undistributed as dividends to the stockholders is termed as retained earnings.

Net income: The net income is the excess of total revenues over total expenses.

Calculate the outstanding common shares.

Particulars
Beginning balance
Less: Treasury stock
Add: Stock dividend
Outstanding common shares
5-Jan
40,000
-
Aprl 5
40.000

Working note:

Calculate stock dividend for October 4.

Stockdividend=((BeginningbalanceofcommonstockTreasurystock)×Percentageofstockdividenddeclared)=(40,0003,000)×20%=37,000×20%=7,400\begin{array}{c}\\{\rm{Stock dividend = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Beginning balance of common stock }}\\\\ - {\rm{ Treasury stock}}\\\end{array} \right){\rm{ }}\\\\ \times {\rm{ Percentage of stock dividend declared}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = }}\left( {{\rm{40,000 }} - {\rm{ 3,000}}} \right){\rm{ }} \times {\rm{ 20\% }}\\\\{\rm{ = 37,000 }} \times {\rm{ 20\% }}\\\\{\rm{ = 7,400}}\\\end{array}

Therefore, the stock dividend for October 4 is 7,400.

Calculate cash dividend amounts.

Particulars
Number of outstanding common shares (a)
Dividend per shares (b)
Cash dividend amounts (a) (b)
5-Jan
40.000
$0.50

Calculate capitalization amount.

Capitalizationamount=Stockdividend×Marketvaluepershare=7,400×$12=$88,800\begin{array}{c}\\{\rm{Capitalization amount = Stock dividend }} \times {\rm{ Market value per share }}\\\\{\rm{ = 7,400}}\; \times \;\$ {\rm{12 }}\\\\{\rm{ = \$ 88,800}}\\\end{array}

Therefore, the capitalization amount is $88,800.

Working note:

Calculate stock dividend for October 4.

Stockdividend=((BeginningbalanceofcommonstockTreasurystock)×Percentageofstockdividenddeclared)=(40,0003,000)×20%=37,000×20%=7,400\begin{array}{c}\\{\rm{Stock dividend = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Beginning balance of common stock }}\\\\ - {\rm{ Treasury stock}}\\\end{array} \right){\rm{ }}\\\\ \times {\rm{ Percentage of stock dividend declared}}\\\end{array} \right){\rm{ }}\\\\{\rm{ = }}\left( {{\rm{40,000 }} - {\rm{ 3,000}}} \right){\rm{ }} \times {\rm{ 20\% }}\\\\{\rm{ = 37,000 }} \times {\rm{ 20\% }}\\\\{\rm{ = 7,400}}\\\end{array}

Calculate per share cost of treasury stock purchased.

Persharecostoftreasurystockpurchased}=CostoftreasurystockNumberofsharesheldastreasury=$30,0003,000=$10\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Per share cost of treasury}}\\\\{\rm{ stock purchased}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{Cost of treasury stock}}}}{{{\rm{Number of shares held as treasury}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 30,000}}}}{{{\rm{3,000}}}}{\rm{ }}\\\\{\rm{ = \$ 10}}\\\end{array}

Therefore, the per share cost of treasury stock purchased is $10.

Calculate net income earned during 2014.

Amount
Amount
$400,000
Particulars
Retained earnings, Dec 31, 2014
Add: Dividend paid
Cash dividend declared on Jan 5
Cash di

Therefore, the net income earned during 2014 is $248,000.

Ans:

5-Jan
40,000
5-Apr
37,000
5-Jul
37,000
5-Oct
44,400
Outstanding common shares

5
Particulars
Cash dividend amounts
5-Jan
$20,000
Aprl
5
$18,500
-Jul
$18,500
5-Oct
$22,200

Capitalization amount is $88,800.

Per share cost of treasury stock purchased is $10.

Net income earned during 2014 is $248,000.